# Edge Universality for Inhomogeneous Random Matrices

### v1

@@ 5 Universality for sub-Gaussian inhomogenous matrices
We need to introduce some necessary notations where the first two ones have been defined in [EK11a, Section 6.1].
@@
where the sum ranges over $l_{i}\geq 0$ for $i=0,\ldots,k$, and $a=(a_{1},\ldots.a_{k})$.
@@
$\left|\mathbb{E}\Big{[}\prod_{j=1}^{s}\mathrm{Tr}\Big{(}V_{l^{(j)}_{0}}\underline{\Phi_{a_{1}^{(j)}}V_{l_{1}^{(j)}}}\cdots\underline{\Phi_{a_{k_{j}}^{(j)}}V_{l_{k_{j}}^{(j)}}}\Big{)}\Big{]}\right|\leq 4^{k_{1}+\cdots+k_{s}}\mathbb{E}\Big{[}\prod_{j=1}^{s}\mathrm{Tr}\Big{(}\widehat{V}_{l^{(j)}_{0}}\underline{\widehat{\Phi_{2}V_{l_{1}^{(j)}}}}\cdots\underline{\widehat{\Phi_{2}V_{l_{k_{j}}^{(j)}}}}\Big{)}\Big{]}$. (5.9)
@@
An immediate corollary from Lemma 5.4 holds.
@@
    Forbidden backtracks. By construction of the nonbacktracking powers $V_{n}$, no two consecutive HH–edges may pair. However, when inserting operators $\underline{\widehat{\Phi_{a_{i}}V_{l_{i}}}}$, those forced “gaps” (the indicator $\mathbf{1}(x_{i}\neq x_{i+2})$) are not imposed across the insertion points, see Figure 8 for example). Consequently, pairings may create local backtracks—these show up as extra “tails” in the diagram. We call such configurations diagrams with tail edges, analogous to cases without red points in [FS10, Sod10, LZ24b].
@@
      there is no restriction for $y$, so $y=x_{1}$ and $y=x_{-1}$ become possible, again creating a tail.
@@
Fix a red point $v\not\in V$ (distinct from all original diagram vertices). The vertex labeling at $v$ admits two configurations in Figure 8: Case(i): $\eta({x})\neq\eta(x_{-2})$ and $\eta(x)\neq\eta(x_{2})$ (Figure 8); Case (ii): $\eta({x})=\eta(x_{-2})$ or $\eta(x)=\eta(x_{2})$ (Figure 10). In case (ii), $x_{-2}x_{-1}x$ forms a backtracking path, and thus contributes a tail to the diagram.
@@
Here the last line we use the Hölder inequality and bounds in Lemma 5.6.
@@
By Corollary 5.8, the mixed moments of Chebyshev polynomials under the conditions of Theorem 1.3 exhibit identical asymptotics to the GOE/GUE cases. The proof of Theorem 1.3 thus follows from almost same procedures as in the Gaussian case. ∎
@@ 6 Extension to inhomogeneous Wishart matrices
$\mathbb{E}\Bigl{[}\prod_{j=1}^{s}\mathrm{Tr}\big{(}X^{m_{j}}\big{)}\Bigr{]}=\sum_{\Gamma\in\mathscr{D}_{s;\beta}}\;\sum_{\begin{subarray}{c}\eta:V(\Gamma)\to[M]\cup[N]\end{subarray}}\;\sum_{\begin{subarray}{c}w_{e}\geq 1\\ \sum_{e\in\partial D_{j}}w_{e}=n_{j},\ \forall j\end{subarray}}\;\Biggl{[}\,\prod_{j=1}^{s}C\bigl{(}m_{j},n_{j}\bigr{)}\;\Biggr{]}\;\\ \times\Biggl{[}\,\prod_{e=(x,y)\in E_{\mathrm{int}}}p_{w_{e}}\bigl{(}\eta(x),\eta(y)\bigr{)}\;\prod_{e=(z,w)\in E_{b}}(A^{(w_{e})})_{\eta(z)\eta(w)}\Biggr{]}.$ (6.14)
@@
we know the Catalan trees contribute the weight of
$\prod_{j=1}^{s}C\bigl{(}m_{j},n_{j}\bigr{)},$ (6.21)
@@
Step 2. It is a well-known fact (e.g. [KMS07, Theorem 1]) that the centered and shifted Chebyshev polynomials
@@
is orthogonal with respect to $\mu_{\mathrm{MP}}$, as established in [FS10, Prop. I.5.2]. Direct moment-matching on $\mu_{\mathrm{MP}}$ consequently yields expansion (6.26). ∎
Define the polynomial
$Q_{n}(x)=U_{n}\!\bigl{(}\tfrac{x-(1+\alpha)}{2\sqrt{\alpha}}\bigr{)}\;+\;\sqrt{\alpha}\,U_{n-1}\!\bigl{(}\tfrac{x-(1+\alpha)}{2\sqrt{\alpha}}\bigr{)}.$ (6.30)
For its path expansion, see Proposition B.4. Then, adapting the argument for inhomogeneous Wigner matrices yields
@@
For $\alpha\in(0,1]$, with $X_{N}=(H_{N}+A_{N})(H_{N}+A_{N})^{*}$ given in Definition 6.1, assume that $A_{N}$ has $q$ critical eigenvalues $a_{j}=\sqrt{\alpha}+\tau_{j}N^{-1/3}$ for $j=1,\ldots,q$ with $\tau_{j}\in\mathbb{R}$, and rqr-q bulk eigenvalues $a_{j}\in(-\sqrt{\alpha},\sqrt{\alpha})$ for $j=q+1,\ldots,r$, where $0\leq q\leq r$ is fixed. If $(P_{S},S)$ satisfies the Short-to-Long Mixing conditions (D1) and (D2) with $\theta t_{N}\ll N^{\frac{1}{3}}$, then the first $k$ largest eigenvalues of $X_{N}$ converge in distribution to those of deformed GOE/GUE matrix. It also holds for first $k$ smallest eigenvalues when $\alpha<1$.
@@ 7.1 Applications
Consider generalized Wigner matrices with $N$-dependent variance lower and upper bounds $c_{N}$ and $C_{N}$. This allows $c_{N}$ and $C_{N}$ to grow or decay with $N$, extending earlier results—see, e.g., [BEY14]—where $c_{N}$ and $C_{N}$ are assumed to be fixed constants.
@@
$\sum_{j}\sigma_{ij}^{2}=1,\qquad c_{N}<N\,\sigma_{ij}^{2}<C_{N},\qquad 0<c_{N}\leq C_{N}<\infty,$ (7.1)
@@
By Theorem 1.3, any generalized Wigner matrix that satisfies (7.1) and has $\theta$-sub-Gaussian entries, with $\theta t_{N}\ll N^{\frac{1}{3}}$, exhibits Tracy-Widom edge universality, provided the constraint (7.2). ∎
@@ 7.1.1 Random band matrices
ensures $\sum_{y}\sigma_{xy}^{2}=1$. The resulting matrix $H$ is a random band matrix of size $N:=L^{d}$ on $d$-dimensional periodic lattice. Applying the local central limit theorem and upper bound for the transition probability, as established in e.g. [GL24], we can easily verify condition (B1) and (B2).
@@
Power-law random band matrices. If the profile $f$ has a finite second moment (so that the small-frequency exponent is $\alpha=2$ in (7.12)), the classical central limit theorem applies and the associated random walk is diffusive with a Gaussian fixed point;
By contrast, when $f$ exhibits heavy tails with decay $|x|^{-d-\alpha}$ for some $0<\alpha<2$, the CLT breaks down and the random walk converges to an $\alpha$-stable process. In that heavy-tailed regime the mixing is faster and the critical bandwidth changes accordingly; the precise power-law thresholds are given in (7.15) and reflect the $\alpha$-stable scaling.
Theorem 1.3 and 7.3 treat both the Gaussian attraction domain ($\alpha=2$) and the stable attraction domain ($0<\alpha<2$) in a unified way.
Non-periodic and structured variants. It is straightforward to check that, under the same assumptions, conditions (B1) and (B2) also hold in the non-periodic setting. However, in order to preserve the doubly stochastic structure of the variance profile, the transition probability should be chosen to correspond to a reflecting random walk. This yields a partial answer to the problem stated in [LZ24b, Section 6, Remark V] in the super-critical regime.
@@ 7.1.2 Random block matrices
Define a block tridiagonal model that is also know as the block Wegner orbital model [Weg79, PSSS19]
@@
Here every $H_{i}$ is an $M\times M$ Wigner matrix with $\theta$-sub-Gaussian entries and every $A_{i}$ is an independent non-Hermitian $M\times M$ matrix with i.i.d. $\theta$-sub-Gaussian entries, and all matrices are independent. This model has been studied in[PSSS19, FSYY25, SYY25, YY25a, KYY25].
An interesting problem is the phase transition for eigenvalue statistics: when $\lambda=0$, the eigenvalue statistics come from $D$ independent matrices; for $\lambda=\frac{1}{2}$, the model is the block random band matrices in dimension $d=1$; for $\lambda=1$, it is block bi-diagonal model.
@@
Identify the block index set with $\mathbb{Z}/D\mathbb{Z}$, and let $q_{n}(a,b)$ be the $n$-step transition probability of the Markov chain that stays at the same block w.p.  $1-\lambda$ and moves to a neighbor w.p.  $\tfrac{\lambda}{2}$. Then for $x$ in block $a$ and $y$ in block $b$,
@@
Similar conclusions hold in higher dimensions, which we omit for brevity. We also note, following [FSYY25], that when $D$ is fixed, a sharp phase transition in the edge statistics occurs at $\|A\|_{\mathrm{HS}}\sim N^{1/3}$, which is equivalent to $\lambda\sim N^{-1/3}$. Therefore, the condition (7.21) in Theorem 7.4 is optimal for fixed $D$. Moreover, in contrast to the settings of [PSSS19, FSYY25, SYY25, YY25a, KYY25], the matrix entries in our model are neither restricted to be Hermitian nor Gaussian, and may even exhibit heavy-tailed behavior and more general variance profile.
@@ 7.1.3 Inhomogeneous sparse random matrices
Previous results [EKYY12, EKYY13b] show that in the homogeneous case, where $w_{ij}=1$ and $p_{ij}=\theta^{-1}$, the largest eigenvalue obeys Tracy-Widom whenever $\theta^{-1}\;\gg\;N^{-1/3}.$ In our framework, one verifies that Assumptions (B1)-(B2) hold with $t_{N}=1$ thus as long as $\theta\;\ll\;N^{1/3},$ Theorem 7.5 yields edge universality in this sparse regime. Moreover, Lee-Schnelli [LS18] proved that for homogeneous Erdős-Rényi graph if $\theta\;\gtrsim\;N^{1/3},$ then a deterministic shift appears at the spectral edge, showing this threshold is sharp. See e.g. [HY22, He23] and references therein for recent results and progress for edge universality of homogeneous sparse Erdős-Rényi graph.
@@ 7.1.4 IRM with heavy-tailed entries
If the IRM matrix $X$ in Definition 1.1 satisfies Assumptions (B1)(B2) with $t_{N}\ll N^{\tfrac{1}{3}-\zeta}$ for some $\zeta\in(0,\tfrac{1}{3})$, and the entries have uniform moments with some fixed $\epsilon_{0}>0$,
@@
Edge universality holds for heavy-tailed Wigner and sample-covariance matrices when entries have only polynomial-decay tails, as shown in [Ruz06, Kho12, LY14, DY18] and references therein.
As a concrete application of Theorem 7.6, for Wigner matrices whose entries satisfy $\mathbb{E}[|W_{ij}|^{12+\epsilon}]<\infty$, edge universality still holds true. This result is somewhat weaker than those in [LY14, DY18]. The distinction lies in the regime of interest: if the distributions of $W_{ij}$ vary with $N$, and their fourth moments diverge as $\mathbb{E}[|W_{ij}|^{4}]\geq N^{\frac{1}{3}}$, then a non-universal correction emerges near the spectral edge (see [LS18] for examples). In contrast, for fixed heavy-tailed distributions, the truncated matrix has much smaller fourth moments, well below the $N^{1/3}$ threshold. For critical tail behavior in random band or sparse matrices, we refer to [Dia23, Han24a].

### v2

@@ 5 Universality for sub-Gaussian inhomogenous matrices
We need to introduce some necessary notations where the first two ones have been defined in [EK11Quantum, Section 6.1].
@@
where the sum ranges over $l_{i}\geq 0$ for $i=0,\ldots,k$, and $a=(a_{1},\ldots,a_{k})$.
@@
[…]
@@
An immediate result of Lemma 5.4 follows.
@@
    Forbidden backtracks. By construction of the nonbacktracking powers $V_{n}$, no two consecutive HH–edges may pair. However, when inserting operators $\underline{\widehat{\Phi_{a_{i}}V_{l_{i}}}}$, those forced “gaps” (the indicator $\mathbf{1}(x_{i}\neq x_{i+2})$) are not imposed across the insertion points, see Figure 8 for example). Consequently, pairings may create local backtracks—these show up as extra “tails” in the diagram. We call such configurations diagrams with tail edges, analogous to cases without red points in [feldheim2010universality, sodin2010spectral, liu2023edge].
@@
      then there is no restriction for $y$, so $y=x_{1}$ and $y=x_{-1}$ become possible, again creating a tail.
@@
Fix a red point $v\not\in V$, distinct from all original diagram vertices. The vertex labeling at $v$ admits two configurations in Figure 8: Case(i): $\eta({x})\neq\eta(x_{-2})$ and $\eta(x)\neq\eta(x_{2})$ (Figure 8); Case (ii): $\eta({x})=\eta(x_{-2})$ or $\eta(x)=\eta(x_{2})$ (Figure 10). In case (ii), $x_{-2}x_{-1}x$ forms a backtracking path, and thus contributes a tail to the diagram.
@@
Here in the last line we use the Hölder inequality and bounds in Lemma 5.6.
@@
Remark 5.1.
The proof of Corollary 5.8 for the deformed case is similar but requires slight modifications. We first adopt the Chebyshev expansion (Proposition A.1) for deformed case and sum over the subscripts of consecutive $\Phi_{A}$ to obtain $(A^{k})_{xy}$ in $\mathcal{D}_{n}$. Then, we apply Lemma 5.4 to establish an upper bound by replacing $(A^{k})_{xy}$ with its absolute value, $|(A^{k})_{xy}|$, similar to the step in Lemma 5.5. From this point, we proceed with the Wick pairing step from Lemma 5.6 to obtain an upper bound for the coupling of the inner edges. Following the argument in Lemma 5.6, the diagram function is consequently bounded by the right-hand side of (3.1) (from the Gaussian case), where the $(A^{k})_{xy}$ terms have been replaced by $|(A^{k})_{xy}|$, and the remainder of the proof proceeds identically.
To conclude this section, we complete the proof of the general case of Theorem 1.3.
@@
By Corollary 5.8, the mixed moments of Chebyshev polynomials under the conditions of Theorem 1.3 exhibit asymptotics identical to those of the GOE/GUE ensemble. The proof of Theorem 1.3 therefore follows a procedure largely identical to the Gaussian case. ∎
@@ 6 Extension to inhomogeneous Wishart matrices
$\mathbb{E}\Bigl[\prod_{j=1}^{s}\mathrm{Tr}\big(X^{m_{j}}\big)\Bigr]=\sum_{\Gamma\in\mathscr{D}_{s;\beta}}\;\sum_{\begin{subarray}{c}\eta:V(\Gamma)\to[M]\cup[N]\end{subarray}}\;\sum_{\begin{subarray}{c}w_{e}\geq 1\\ \sum_{e\in\partial D_{j}}w_{e}=n_{j},\ \forall j\end{subarray}}\;\Biggl[\,\prod_{j=1}^{s}C_{\mathrm{MP}}(m_{j},n_{j})\;\Biggr]\;\\ \times\Biggl[\,\prod_{e=(x,y)\in E_{\mathrm{int}}}p_{w_{e}}\bigl(\eta(x),\eta(y)\bigr)\;\prod_{e=(z,w)\in E_{b}}(A^{(w_{e})})_{\eta(z)\eta(w)}\Biggr].$ (6.14)
@@
we know that the Catalan trees contribute the weight of
$\prod_{j=1}^{s}C_{\mathrm{MP}}(m_{j},n_{j}),$ (6.21)
@@
Step 2. It is a well-known fact (e.g. [MR2320312, Theorem 1]) that the centered and shifted Chebyshev polynomials
@@
is orthogonal with respect to $\mu_{\mathrm{MP}}$, as established in [feldheim2010universality, Prop. I.5.2]. Direct moment-matching on $\mu_{\mathrm{MP}}$ consequently yields expansion (6.26). ∎
Introduce a family of polynomials
$Q_{n}(x)=U_{n}\!\bigl(\tfrac{x-(1+\alpha)}{2\sqrt{\alpha}}\bigr)\;+\;\sqrt{\alpha}\,U_{n-1}\!\bigl(\tfrac{x-(1+\alpha)}{2\sqrt{\alpha}}\bigr),$ (6.30)
then we can obtain a path expansion; see Proposition B.4. Then, adapting the argument for inhomogeneous Wigner matrices yields
@@
For $\alpha\in(0,1]$, with $X_{N}=(H_{N}+A_{N})(H_{N}+A_{N})^{*}$ given in Definition 6.1, assume that $A_{N}$ has $q$ critical eigenvalues $a_{j}=\sqrt{\alpha}+\tau_{j}N^{-1/3}$ for $j=1,\ldots,q$ with $\tau_{j}\in\mathbb{R}$, and rqr-q bulk eigenvalues $a_{j}\in(-\sqrt{\alpha},\sqrt{\alpha})$ for $j=q+1,\ldots,r$, where $0\leq q\leq r$ is fixed. If $(P_{S},S)$ satisfies the Short-to-Long Mixing conditions (D1) and (D2) with $\theta t_{N}\ll N^{\frac{1}{3}}$, then the first $k$ largest eigenvalues of $X_{N}$ converge in distribution to those of deformed GOE/GUE matrix. Similar results hold for first $k$ smallest eigenvalues when $\alpha<1$.
@@
Theorem 6.7 is broadly applicable, much like Theorem 1.3 (see Section 7.1 below). These include non-Hermitian matrices $H_{N}$ in (6.1) chosen as random band matrices with independent entries and general variance profiles, as well as sparse random matrices with structured variance profiles.
@@ 7.1 Applications
7.1.1 Generalized Wigner matrices
Consider generalized Wigner matrices with $N$-dependent variance lower and upper bounds $c_{N}$ and $C_{N}$. This allows $c_{N}$ and $C_{N}$ to grow or decay with $N$, extending earlier results—see, e.g., [MR3253704]—where $c_{N}$ and $C_{N}$ are assumed to be fixed constants.
@@ 7.1.1 Generalized Wigner matrices
$\sum_{j}\sigma_{ij}^{2}=1,\quad 0<c_{N}<N\,\sigma_{ij}^{2}<C_{N}<\infty,$ (7.1)
@@
By Theorem 1.3, any generalized Wigner matrix that satisfies (7.1) and has $\theta$-sub-Gaussian entries, with $\theta t_{N}\ll N^{{1}/{3}}$, exhibits Tracy-Widom edge universality, provided the constraint (7.2). ∎
@@ 7.1.2 Random band matrices
ensures $\sum_{y}\sigma_{xy}^{2}=1$. The resulting matrix $H$ is a random band matrix of size $N:=L^{d}$ on $d$-dimensional periodic lattice. Applying the local central limit theorem and upper bound for the transition probability, as established in e.g. [gu2024local], we can easily verify condition (B1) and (B2).
@@
Power-law random band matrices. If the profile $f$ has a finite second moment (so that the small-frequency exponent is $\alpha=2$ in (7.12)), the classical central limit theorem applies and the associated random walk is diffusive with a Gaussian fixed point. By contrast, when $f$ exhibits heavy tails with decay $|x|^{-d-\alpha}$ for some $0<\alpha<2$, the CLT breaks down and the random walk converges to an $\alpha$-stable process. In that heavy-tailed regime the mixing is faster and the critical bandwidth changes accordingly; the precise power-law thresholds are given in (7.15) and reflect the $\alpha$-stable scaling. Theorems 1.3 and 7.3 treat both the Gaussian attraction domain ($\alpha=2$) and the stable attraction domain ($0<\alpha<2$) in a unified way.
Non-periodic and structured variants. It is straightforward to check that, under the same assumptions, conditions (B1) and (B2) also hold in the non-periodic setting. However, in order to preserve the doubly stochastic structure of the variance profile, the transition probability should be chosen to correspond to a reflecting random walk. This yields a partial answer to the problem stated in [liu2023edge, Section 6, Remark V] in the super-critical regime.
@@ 7.1.3 Random block matrices
Define a block tridiagonal model that is also know as the block Wegner orbital model [wegner1979disordered, MR3915294]
@@
Here every $H_{i}$ is an $M\times M$ Wigner matrix with $\theta$-sub-Gaussian entries and every $A_{i}$ is an independent non-Hermitian $M\times M$ matrix with i.i.d. $\theta$-sub-Gaussian entries, and all matrices are independent. This model has been studied in[MR3915294, fan2025localization, stone2025random, yang2025delocalization, khang2025localization].
An interesting problem is the phase transition for eigenvalue statistics: when $\lambda=0$, the eigenvalue statistics come from $D$ independent matrices; for $\lambda={1}/{2}$, the model is a block random band matrix in dimension $d=1$; for $\lambda=1$, it is a block bi-diagonal model.
@@
Identify the block index set with $\mathbb{Z}/D\mathbb{Z}$, and let $q_{n}(a,b)$ be the $n$-step transition probability of the Markov chain that stays at the same block with probability $1-\lambda$ and moves to a neighbor with probability $\lambda/2$. Then for $x$ in block $a$ and $y$ in block $b$,
@@
Similar conclusions hold in higher dimensions, which we omit for brevity. We also note, following [fan2025localization], that when $D$ is fixed, a sharp phase transition in the edge statistics occurs at $\|A\|_{\mathrm{HS}}\sim N^{1/3}$, which is equivalent to $\lambda\sim N^{-1/3}$. Therefore, the condition (7.21) in Theorem 7.4 is optimal for fixed $D$. Moreover, in contrast to the settings of [MR3915294, fan2025localization, stone2025random, yang2025delocalization, khang2025localization], the matrix entries in our model are neither restricted to be Hermitian nor Gaussian, and may even exhibit heavy-tailed behavior and more general variance profile.
@@ 7.1.4 Inhomogeneous sparse random matrices
Previous results [MR2964770, MR3098073] show that in the homogeneous case, where $w_{ij}=1$ and $p_{ij}=\theta^{-1}$, the largest eigenvalue obeys Tracy-Widom whenever $\theta^{-1}\;\gg\;N^{-1/3}.$ In our framework, one verifies that Assumptions (B1)-(B2) hold with $t_{N}=1$ thus as long as $\theta\;\ll\;N^{1/3},$ Theorem 7.5 yields edge universality in this sparse regime. Moreover, Lee-Schnelli [lee2018] proved that for homogeneous Erdős-Rényi graph if $\theta\;\gtrsim\;N^{1/3},$ then a deterministic shift appears at the spectral edge, showing this threshold is sharp. See e.g. [huang2022edge, he2023edge] and references therein for recent results and progress for edge universality of homogeneous sparse Erdős-Rényi graph.
@@ 7.1.5 IRM with heavy-tailed entries
If the IRM matrix $X$ in Definition 1.1 satisfies Assumptions (B1)(B2) with $t_{N}\ll N^{\tfrac{1}{3}-\zeta}$ for some $\zeta\in(0,\tfrac{1}{3})$, and the entries of $W$ have uniform moments with some fixed $\epsilon_{0}>0$,
@@
Edge universality holds for heavy-tailed Wigner and sample-covariance matrices when entries have only polynomial-decay tails, as shown in [MR2191882, MR2899796, MR3161313, MR3809475] and references therein.
As a concrete application of Theorem 7.6, for Wigner matrices whose entries satisfy $\mathbb{E}[|W_{ij}|^{12+\epsilon}]<\infty$, edge universality still holds true. This result is somewhat weaker than those in [MR3161313, MR3809475]. The distinction lies in the regime of interest: if the distributions of $W_{ij}$ vary with $N$, and their fourth moments diverge as $\mathbb{E}[|W_{ij}|^{4}]\geq N^{\frac{1}{3}}$, then a non-universal correction emerges near the spectral edge (see [lee2018] for examples). In contrast, for fixed heavy-tailed distributions, the truncated matrix has much smaller fourth moments, well below the $N^{1/3}$ threshold. For critical tail behavior in random band or sparse matrices, we refer to [MR4546632, han2024deformed].